# The Behrens–Fisher problem revisited

### v1

@@ Abstract
We revisit the two-sample Behrens–Fisher problem—testing equality of means when two normal populations have unequal, unknown variances—and derive a compact expression for the null distribution of the classical test statistic. The key step is a Mellin–Barnes factorization that decouples the square root of a weighted sum of independent chi-square variates, thereby collapsing a challenging two-dimensional integral to a tractable single-contour integral. Closing the contour yields a residue series that terminates whenever either sample’s degrees of freedom is odd. A complementary Euler–Beta reduction identifies the density as a Gauss hypergeometric function with explicit parameters, yielding a numerically stable form that recovers Student’s $t$ under equal variances. Ramanujan’s master theorem supplies exact inverse-power tail coefficients, which bound Lugannani–Rice saddle-point approximation errors and support reliable tail analyses. Our result subsumes the hypergeometric density derived by Nel et al., and extends it with a concise cdf and analytic tail expansions; their algebraic special cases coincide with our truncated residue series. Using our derived expressions, we tabulate exact two-sided critical values over a broad grid of sample sizes and variance ratios that reveal the parameter surface on which the well-known Welch’s approximation switches from conservative to liberal, quantifying its maximum size distortion.
Keywords: Mellin–Barnes integral; hypergeometric functions; Ramanujan’s theorem.
@@ 1 Introduction
Consider independent samples $X_{1},\dots,X_{n_{1}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ and $Y_{1},\dots,Y_{n_{2}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ from two normal populations. Let $\nu_{i}=n_{i}-1$, $\bar{X}=\tfrac{1}{n_{1}}\sum_{i=1}^{n_{1}}X_{i}$, $\bar{Y}=\tfrac{1}{n_{2}}\sum_{i=1}^{n_{2}}Y_{i}$, $S_{1}^{2}=\tfrac{1}{\nu_{1}}\sum_{i=1}^{n_{1}}(X_{i}-\bar{X})^{2}$, $S_{2}^{2}=\tfrac{1}{\nu_{2}}\sum_{i=1}^{n_{2}}(Y_{i}-\bar{Y})^{2}$ denote the degrees of freedom, the sample means and the sample variances, respectively, of the two samples. For notational convenience, let $g=\sigma_{1}^{2}/n_{1}$, $h=\sigma_{2}^{2}/n_{2}$ and variance ratio $r=\sigma_{1}^{2}/\sigma_{2}^{2}$. Following the pioneering work of Behrens [1] and Fisher [2], the problem of testing $H_{0}\!:\mu_{1}=\mu_{2}$ with $(\sigma_{1}^{2},\sigma_{2}^{2})$ unknown and $\sigma_{1}^{2}\neq\sigma_{2}^{2}$ (i.e., $g\neq h$) remains an enduring challenge in mathematical statistics. Although several approximations, modifications and practical procedures have been proposed (for example, [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15])—including nonparametric versions of the test [16, 17, 18, 19]—to the best of our knowledge, no explicit canonical expression for the density function of the Behrens–Fisher statistic
@@
has been established in terms of standard probability distribution functions. The absence of such a representation complicates exact significance evaluation, power calculations, and asymptotic error control—particularly in the small-sample, high-variance-imbalance regimes common in modern experimental design.
@@
$T$ $=$ $\frac{\bar{X}-\bar{Y}}{\sqrt{\frac{gW_{1}}{\nu_{1}}+\frac{hW_{2}}{\nu_{2}}}},$ (2)
@@
$T\left|\sigma_{1}^{2},\sigma_{2}^{2}\right.\sim\frac{\mathcal{N}\left(0,g+h\right)}{\sqrt{\frac{gW_{1}}{\nu_{1}}+\frac{hW_{2}}{\nu_{2}}}}.$ (3)
@@
$T\left|\sigma^{2}\right.\sim\frac{\left(\bar{X}-\bar{Y}\right)}{\sqrt{S^{2}_{p}\left(\frac{1}{n_{1}}+\frac{1}{n_{2}}\right)}},$ (4)
@@
$S^{2}_{p}=\frac{1}{\nu_{1}+\nu_{2}}\left(\sum_{j=1}^{n_{1}}(X_{j}-\bar{X})^{2}+\sum_{j=1}^{n_{2}}(Y_{j}-\bar{Y})^{2}\right).$ (5)
This implies $T\left|\sigma^{2}\right.\sim t_{\nu_{1}+\nu_{2}}$, the Student’s two-sample $t$-test following a $t$-distribution with $(\nu_{1}+\nu_{2})$ degrees of freedom (see [20, Exercise 8.41]). However, when $\sigma_{1}^{2}\neq\sigma_{2}^{2}$ and $(\sigma_{1}^{2},\sigma_{2}^{2})$ unknown, no such simplification—including using the pooled variance estimate $S^{2}_{p}$—is possible (see [20, Exercise 8.42]). The distribution of
$T$ $=$ $\frac{\bar{X}-\bar{Y}}{\sqrt{\frac{gW_{1}}{\nu_{1}}+\frac{hW_{2}}{\nu_{2}}}}$ (6)
would require integrating over the joint density of $(W_{1},W_{2})$:
$f_{T}(t)=\int_{0}^{\infty}\int_{0}^{\infty}f_{T|W_{1},W_{2}}(t)f_{W_{1}}(w_{1})f_{W_{2}}(w_{2})\,dw_{1}\,dw_{2},$ (7)
@@
Efforts to solve (8) does not yield a closed-form expression in terms of elementary distributions. This complexity is the centerpiece of the Behrens–Fisher problem: lifting the equal-variance assumption transforms the otherwise straightforward two-sample $t$-test into a setting where no exact, closed-form distribution is available in elementary terms. In addressing this problem, Welch [21] employed the approximation developed by Satterthwaite [22] to replace the unknown degrees of freedom with a carefully crafted estimate,
$\hat{\nu}\approx\frac{\left(\frac{S_{1}^{2}}{n_{1}}+\frac{S_{2}^{2}}{n_{2}}\right)^{2}}{\left[\frac{S_{1}^{4}}{n^{2}_{1}\nu_{1}}+\frac{S_{2}^{4}}{n^{2}_{2}\nu_{2}}\right]},$ (9)
@@
thereby yielding $T\overset{\text{approx}}{\sim}t_{\hat{\nu}}$ with $\hat{\nu}$ degrees of freedom. Welch’s test controls the Type I error rate, does not require equal variances and is more robust than the classic pooled-variance $t$-test (see [20, Chapter 8]). The Welch–Satterthwaite approximation has been the main workhorse for the Behrens–Fisher problem in practical applications.
@@
We make progress on this problem by deriving an expression for the null distribution of $T$ that is simultaneously compact, rapidly computable and amenable to asymptotic analyses. Leveraging a Mellin–Barnes (MB) contour decomposition [23], we disentangle the square-root coupling $(gW_{1}/\nu_{1}+hW_{2}/\nu_{2})^{1/2}$ between the independent $\chi^{2}$-variates $W_{1}$ and $W_{2}$. This step collapses the conventional two-dimensional integral representation of the density into a one-dimensional contour integral whose integrand is a product of Gamma functions. Such MB collapses have proved potent in analytic number theory and special-function theory, yet have not previously been applied to the Behrens–Fisher problem. Closing the MB contour on the right yields a hypergeometric-type residue series. The infinite sum thus obtained becomes a finite polynomial-over-rational expression—no tail terms of order $k>\tfrac{\nu_{i}-1}{2}$ survive. The series involves the factor $\Gamma\left(\tfrac{\nu_{2}}{2}-k+\tfrac{1}{2}\right)$, $k=0,1,\dots$, and if, say, $\nu_{2}=2m+1$ (i.e., odd) then the argument hits a non-positive integer at $k=m$; the Gamma function is infinite there, but its reciprocal appears in the residue coefficient, giving zero residue for every $k>m$. Hence, only the first $m+1$ terms remain and the sum terminates. An analogous cut-off occurs when $\nu_{1}$ is odd. This finite-sum phenomenon explains scattered algebraic densities reported in numerical tables from the 1950s (see, for instance, [24, Table 55]) but never formally derived; our terminating series supplies the missing derivation and shows that every odd $\nu_{i}$ row in the Biometrika tables is the closed residue sum, not a separate special case. For general $\nu_{1},\nu_{2}$, the resulting infinite series converges at a geometric rate controlled by $\min\{g,h\}/(g+h)$, enabling high-precision density evaluations with fewer than ten terms across virtually the entire support.
We complement the residue expansion by an Euler $(S,U)$-change of variables that separates a radial component $S=gW_{1}/\nu_{1}+hW_{2}/\nu_{2}$ and a Beta-distributed mixing fraction $U$. The integral in $U$ can be performed in a closed form, producing the compact density
$f_{T}(t)$ $=$ $C\left[2/B(t)\right]^{c}{}_{2}F_{1}\!\left(c,\tfrac{\nu_{1}}{2};\tfrac{\nu_{1}+\nu_{2}}{2};-\rho(t)\right),$ (10)
@@
with explicit expressions for $C$ and $B(t)$, where $c=(\nu_{1}+\nu_{2}+1)/2$ and $\rho(t)=(g^{-1}-h^{-1})/\left(h^{-1}+t^{2}/(g+h)\right)$ and $_{2}F_{1}$ is Gauss’ hypergeometric function and all parameters are explicit. The expression collapses seamlessly to the Student-$t_{\nu_{1}+\nu_{2}-2}$ density when $g=h$ and to a rational function when one variance dominates, integrating disparate corner cases into a coherent framework. While hypergeometric representations have appeared in isolated special–case studies (for instance, [11]), our result is valid for all $(n_{1},n_{2},r)$. Its derivation also clarifies structural symmetries—particularly the invariance under $(\nu_{1},\nu_{2},g,h)\mapsto(\nu_{2},\nu_{1},h,g)$—and provides analytic gradients for likelihood-based estimation of variance ratios in random-effect meta-analyses.
@@
The MB contour representation admits two analytically complementary pathways. First, Euler–Beta reduction: by rewriting $W_{1},W_{2}$ in polar coordinates $S=gW_{1}/\nu_{1}+hW_{2}/\nu_{2},U=gW_{1}//\nu_{1}S$, the inner $S$-integral is a Gamma integral and the remaining $U$-integral equals Euler’s Beta integral representation of Gauss’ hypergeometric function [25]. This gives a closed form expression for $f_{T}(t)$ described in the aforementioned paragraph; the same identity collapses to Studentt-t when $g=h$. Second, Ramanujan’s master theorem (RMT) [26] transforms the MB integrand into an inverse–Mellin power series, producing an exact tail expansion $f_{T}(t)=\sum_{m\geq 0}(-1)^{m}A_{m}|t|^{-(2m+1)}/m!$ with $A_{m}$ as closed Gamma products. The remainder after $M$ terms is $O(\lvert t\rvert^{-(2M+3)})$, leading to deterministic error bounds for any truncated series and gives algebraic control unavailable in simulation-based methods. The same cumulant generator feeds the Lugannani–Rice (LR) saddle-point formula $P(T>t)\approx\Phi(-w)+\phi(w)(1/w-1/u)$ [27], with $w$, $u$ expressed in cumulants of $T$. Using RMT coefficients, we prove a uniform remainder $O(\lvert t\rvert^{-4})$ when $\lvert t\rvert<0.8\min\{\nu_{1},\nu_{2}\}$ and quantify the accuracy loss in the variance–extreme corners ($\lvert\log_{10}r\rvert>1$). These three ingredients—Euler–Beta for the bulk, RMT for the far tail, and LR to bridge the regimes—enable fast and accurate evaluation of the density, distribution function and quantiles through closed-form and numerically stable methods, without resort to computationally intensive simulation.
@@
It was first recognized by Nel et al. [28] that the Behrens–Fisher density can be written as a single Gauss hypergeometric function, but they derived the expression only for the univariate case, gave no proof of termination when a degree of freedom is odd, provided neither cumulative distribution function (cdf) nor tail theory, and concluded that their formula was computationally impractical. Our MB treatment reproduces their pdf as one branch of a richer structure: the residue series shows why the hypergeometric terminates when $\nu_{1}$ or $\nu_{2}$ is odd, the Euler–Beta route supplies a closed-form cdf (an Appell $F_{1}$ integral) for the first time, and Ramanujan’s coefficients turn the putative obstacle of numerical instability into an explicit $\lvert t\rvert^{-(2m+1)}$ tail expansion with deterministic error bounds. On the approximation side, Welch’s $t_{\hat{\nu}}$ statistic replaces the true tail exponent $\nu_{1}+\nu_{2}+1$ by a random $\hat{\nu}$; using our exact density we compute critical-value grids that locate the surface where $E[\hat{\nu}]=\nu_{1}+\nu_{2}$ and show that, away from this surface, Welch can be up to 60% liberal or conservative in samples as small as $n_{1}=n_{2}=10$. It is also of interest to note that, [9] constructed exact tests, but only supply critical constants—not a symbolic density. Their method does not reveal the algebraic structure (hypergeometric kernel, termination when $\nu_{i}$ is odd, Ramanujan tails) that our MB analysis uncovers. Their exactness is achieved by numerical double integration for every $(n_{1},n_{2},r)$. We give an expression that works for all $(n_{1},n_{2},r)$ in a single expression.
@@
The remainder of the paper is organized as follows. In Section 2, the Mellin–Barnes representation of $T$ is presented. The density function $f_{T}(t)$ is expressed as a Gauss hypergeometric function $_{2}F_{1}$ by reducing the 2-D integral to a 1-D integral in Section 3. The tail analyses of $f_{T}(t)$ using Ramanujan’s master theorem is in Section 4. Some discussions including cdf tables, Lugannani–Rice saddle-points, and comparisons between our work with Welch’s [21] and Nel et al. [28] are provided in Section 5. Section 6 concludes the paper.

### v2

# A unified approach to the Behrens–Fisher problem

@@ Abstract
A unified framework is presented to study the two-sample Behrens–Fisher problem—testing equality of means when two normal populations have unequal, unknown variances—and a compact expression is derived for the null distribution of the classical test statistic. Our new approach involves a Mellin–Barnes factorization that decouples the square root of a weighted sum of independent chi-square variates, thereby collapsing a challenging two-dimensional integral to a tractable single-contour integral. Closing the contour yields a residue series that terminates whenever either sample’s degrees of freedom is odd. A complementary Euler–Beta reduction identifies the density as a Gauss hypergeometric function with explicit parameters, yielding a numerically stable form that recovers Student’s $t$ under equal variances. Ramanujan’s master theorem supplies exact inverse-power tail coefficients, which bound Lugannani–Rice saddle-point approximation errors and support reliable tail analyses. The proposed framework reveals why hypergeometric structure appears, why certain finite-sum cases arise, and how one can pass from the bulk of the distribution to its tails without altering the analytic framework. Finally, it lets us tabulate exact two-sided critical values over a broad grid of sample sizes and variance ratios that reveal the parameter surface on which the well-known Welch’s approximation switches from conservative to liberal, quantifying its maximum size distortion.
Keywords: Mellin–Barnes integral, hypergeometric functions, Ramanujan’s master theorem.
@@ 1 Introduction
Consider independent samples $X_{1},\dots,X_{n_{1}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ and $Y_{1},\dots,Y_{n_{2}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ from two normal populations. Let $\nu_{i}=n_{i}-1$, $\bar{X}=\tfrac{1}{n_{1}}\sum_{i=1}^{n_{1}}X_{i}$, $\bar{Y}=\tfrac{1}{n_{2}}\sum_{i=1}^{n_{2}}Y_{i}$, $S_{1}^{2}=\tfrac{1}{\nu_{1}}\sum_{i=1}^{n_{1}}(X_{i}-\bar{X})^{2}$, $S_{2}^{2}=\tfrac{1}{\nu_{2}}\sum_{i=1}^{n_{2}}(Y_{i}-\bar{Y})^{2}$ denote the degrees of freedom, the sample means and the sample variances, respectively, of the two samples. Also, let $g=\sigma_{1}^{2}/n_{1}$, $h=\sigma_{2}^{2}/n_{2}$ and variance ratio $r=\sigma_{1}^{2}/\sigma_{2}^{2}$. Following the pioneering work of Behrens (1929) and Fisher (1935), the problem of testing $H_{0}\!:\mu_{1}=\mu_{2}$ with $(\sigma_{1}^{2},\sigma_{2}^{2})$ unknown and $\sigma_{1}^{2}\neq\sigma_{2}^{2}$ remains an enduring challenge in mathematical statistics. Several approximations, modifications and practical procedures have been proposed to address this long-standing problem—for instance, see the work of Scheffé (1970); Bozdogan and Ramirez (1986); Best and Rayner (1987); Asiribo and Gurland (1989); Duong and Shorrock (1996); Kim and Cohen (1998); Vangel (2005); Dudewicz et al. (2007); Chang and Pal (2008); Nadarajah and Li (2017); Chaturvedi et al. (2019); Wang and Xu (2022); Chen et al. (2022, 2023). Other variants including nonparametric versions of the test (see Brunner et al. (2002); Larocque et al. (2010); Konietschke and Pauly (2012); Konietschke et al. (2012); He et al. (2022)) and high-dimensional settings (see, for example, Zhou et al. (2017); Cao et al. (2019); Zhang et al. (2021); Pei et al. (2026)) have also been reported in the literature. Despite these useful and impactful contributions, to the best of our knowledge, no explicit canonical expression for the density function of the Behrens–Fisher statistic
@@
has been established in terms of standard probability distribution functions. The absence of such a representation complicates exact significance evaluation, power calculations, and asymptotic error control—particularly in the small-sample, high-variance-imbalance regimes common in modern experimental design. In what follows, we first explain why the Behrens–Fisher problem remains a persistent difficulty in mathematical statistics, and then in Section 1.1 outline how our work advances its analysis.
@@
$T$ $=$ $\frac{\bar{X}-\bar{Y}}{\sqrt{\frac{gW_{1}}{\nu_{1}}+\frac{hW_{2}}{\nu_{2}}}},$
@@
$T\left|\sigma_{1}^{2},\sigma_{2}^{2}\right.\sim\frac{\mathcal{N}\left(0,g+h\right)}{\sqrt{\frac{gW_{1}}{\nu_{1}}+\frac{hW_{2}}{\nu_{2}}}}.$
@@
$T\left|\sigma^{2}\right.\sim\frac{\left(\bar{X}-\bar{Y}\right)}{\sqrt{S^{2}_{p}\left(\frac{1}{n_{1}}+\frac{1}{n_{2}}\right)}},$
@@
$S^{2}_{p}=\frac{1}{\nu_{1}+\nu_{2}}\left(\sum_{j=1}^{n_{1}}(X_{j}-\bar{X})^{2}+\sum_{j=1}^{n_{2}}(Y_{j}-\bar{Y})^{2}\right).$
This implies $T\left|\sigma^{2}\right.\sim t_{\nu_{1}+\nu_{2}}$, the Student’s two-sample $t$-test following a $t$-distribution with $(\nu_{1}+\nu_{2})$ degrees of freedom (see (Casella and Berger, 2001, Exercise 8.41)). However, when $\sigma_{1}^{2}\neq\sigma_{2}^{2}$ and $(\sigma_{1}^{2},\sigma_{2}^{2})$ unknown, no such simplification—including using the pooled variance estimate $S^{2}_{p}$—is possible (see (Casella and Berger, 2001, Exercise 8.42)). The distribution of $T=\frac{\bar{X}-\bar{Y}}{\sqrt{\frac{gW_{1}}{\nu_{1}}+\frac{hW_{2}}{\nu_{2}}}}$ would require integrating over the joint density of $(W_{1},W_{2})$:
$f_{T}(t)=\int_{0}^{\infty}\int_{0}^{\infty}f_{T|W_{1},W_{2}}(t)f_{W_{1}}(w_{1})f_{W_{2}}(w_{2})\,dw_{1}\,dw_{2},$
@@
Efforts to solve (2) does not yield a closed-form expression in terms of elementary distributions. This complexity is the centerpiece of the Behrens–Fisher problem: lifting the equal-variance assumption transforms the otherwise straightforward two-sample $t$-test into a setting where no exact, closed-form distribution is available in elementary terms. In addressing this problem, Welch (1947) employed the approximation developed by Satterthwaite (1946) to replace the unknown degrees of freedom with a carefully crafted estimate,
$\hat{\nu}\approx\frac{\left(\frac{S_{1}^{2}}{n_{1}}+\frac{S_{2}^{2}}{n_{2}}\right)^{2}}{\left[\frac{S_{1}^{4}}{n^{2}_{1}\nu_{1}}+\frac{S_{2}^{4}}{n^{2}_{2}\nu_{2}}\right]},$
@@
thereby yielding $T\overset{\text{approx}}{\sim}t_{\hat{\nu}}$ with $\hat{\nu}$ degrees of freedom. Welch’s test controls the Type I error rate, does not require equal variances and is more robust than the classic pooled-variance $t$-test ((Casella and Berger, 2001, Chapter 8)). The Welch–Satterthwaite approximation has been the main workhorse for the Behrens–Fisher problem in practical applications.
1.1 Our contribution
@@
We make progress on this problem by deriving an expression for the null distribution of $T$ that is simultaneously compact, rapidly computable and amenable to asymptotic analyses. We propose a new, unified approach that leverages a Mellin–Barnes (MB) contour decomposition (see Paris and Kaminski (2001)), thereby disentangling the square-root coupling $(gW_{1}/\nu_{1}+hW_{2}/\nu_{2})^{1/2}$ between the independent $\chi^{2}$-variates $W_{1}$ and $W_{2}$. This step collapses the conventional two-dimensional integral representation of the density given by (2) into a one-dimensional contour integral whose integrand is a product of Gamma functions. Such MB collapses have proved potent in the analysis of special-function theory, yet they have not previously been applied to the Behrens–Fisher problem, which has been extensively studied from the standpoint of special functions. Closing the MB contour on the right yields a hypergeometric-type residue series. The infinite sum thus obtained becomes a finite polynomial-over-rational expression—no tail terms of order $k>\tfrac{\nu_{i}-1}{2}$ survive. The series involves the factor $\Gamma\left(\tfrac{\nu_{2}}{2}-k+\tfrac{1}{2}\right)$, $k=0,1,\dots$, and if, say, $\nu_{2}=2m+1$ (i.e., odd) then the argument hits a non-positive integer at $k=m$; the Gamma function is infinite there, but its reciprocal appears in the residue coefficient, giving zero residue for every $k>m$. Hence, only the first $m+1$ terms remain and the sum terminates. An analogous cut-off occurs when $\nu_{1}$ is odd. This finite-sum phenomenon explains scattered algebraic densities reported in numerical tables from the 1950s (see, for instance, (Pearson and Hartley, 1958, Table 55)) but never formally derived; our terminating series supplies the missing derivation and shows that every odd $\nu_{i}$ row in the Biometrika tables is the closed residue sum, not a separate special case. For general $\nu_{1},\nu_{2}$, the resulting infinite series converges at a geometric rate controlled by $\min\{g,h\}/(g+h)$, enabling high-precision density evaluations with fewer than ten terms across the entire support.
@@
We complement the residue expansion by an Euler $(S,U)$-change of variables that separates a radial component $S=gW_{1}/\nu_{1}+hW_{2}/\nu_{2}$ and a Beta-distributed mixing fraction $U$. The integral in $U$ can be performed in a closed form, producing the compact density $f_{T}(t)=C\left[2/B(t)\right]^{c}{}_{2}F_{1}\!\left(c,\tfrac{\nu_{1}}{2};\tfrac{\nu_{1}+\nu_{2}}{2};-\rho(t)\right)$, with explicit expressions for $C$ and $B(t)$, where $c=(\nu_{1}+\nu_{2}+1)/2$ and $\rho(t)=(g^{-1}-h^{-1})/\left(h^{-1}+t^{2}/(g+h)\right)$ and $_{2}F_{1}$ is Gauss’ hypergeometric function and all parameters are explicit. The expression collapses seamlessly to the Student-$t_{\nu_{1}+\nu_{2}}$ density when $\sigma_{1}^{2}=\sigma_{2}^{2}$ and to a rational function when one variance dominates, integrating disparate corner cases into a coherent framework. While hypergeometric representations have appeared in isolated special–case studies (for instance, Nadarajah and Li (2017)), our result is valid for all $(n_{1},n_{2},r)$. Its derivation also clarifies structural symmetries—particularly the invariance under $(\nu_{1},\nu_{2},g,h)\mapsto(\nu_{2},\nu_{1},h,g)$—and provides analytic gradients for likelihood-based estimation of variance ratios in random-effect meta-analyses.
@@
The MB contour representation admits two analytically complementary pathways. First, Euler–Beta reduction: by rewriting $W_{1},W_{2}$ in polar coordinates $S=gW_{1}/\nu_{1}+hW_{2}/\nu_{2},U=gW_{1}/\nu_{1}S$, the inner $S$-integral is a Gamma integral and the remaining $U$-integral equals Euler’s Beta integral representation of Gauss’ hypergeometric function (see Andrews et al. (1999)). This gives a closed form expression for $f_{T}(t)$ described in the aforementioned paragraph; the same identity collapses to Studentt-t when $g=h$. Second, Ramanujan’s master theorem (RMT) summarized by Berndt (1985) transforms the MB integrand into an inverse–Mellin power series, producing an exact tail expansion $f_{T}(t)=\sum_{m\geq 0}(-1)^{m}A_{m}|t|^{-(2m+1)}/m!$ with $A_{m}$ as closed Gamma products. The remainder after $M$ terms is $O(\lvert t\rvert^{-(2M+3)})$, leading to deterministic error bounds for any truncated series and gives algebraic control unavailable in simulation-based methods. The same cumulant generator feeds the Lugannani–Rice (LR) saddle-point formula $P(T>t)\approx\Phi(-w)+\phi(w)(1/w-1/u)$ (see Lugannani and Rice (1980)), with $w$, $u$ expressed in cumulants of $T$. Using RMT coefficients, we prove a uniform remainder $O(\lvert t\rvert^{-4})$ when $\lvert t\rvert<0.8\min\{\nu_{1},\nu_{2}\}$ and quantify the accuracy loss in the variance–extreme corners ($\lvert\log_{10}r\rvert>1$). These three ingredients—Euler–Beta for the central region (i.e., bulk), RMT for the far tail, and LR to bridge the regimes—enable fast and accurate evaluation of the density, distribution function and quantiles through closed-form and numerically stable methods, without resort to computationally intensive simulation.
@@
It was first recognized by Nel et al. (1990) that the Behrens–Fisher density can be written as a single Gauss hypergeometric function, but they derived the expression only for the univariate case, gave no proof of termination when a degree of freedom is odd, provided neither cumulative distribution function (cdf) nor tail theory, and concluded that their formula was computationally impractical. Our MB treatment reproduces their pdf as one branch of a richer structure: the residue series shows why the hypergeometric terminates when $\nu_{1}$ or $\nu_{2}$ is odd, the Euler–Beta route supplies a closed-form cdf (an Appell $F_{1}$ integral) for the first time, and Ramanujan’s coefficients turn the putative obstacle of numerical instability into an explicit $\lvert t\rvert^{-(2m+1)}$ tail expansion with deterministic error bounds. On the approximation side, Welch’s $t_{\hat{\nu}}$ statistic replaces the true tail exponent $\nu_{1}+\nu_{2}+1$ by a random $\hat{\nu}$; using our exact density we compute critical-value grids that locate the surface where $E[\hat{\nu}]=\nu_{1}+\nu_{2}$ and show that, away from this surface, Welch can be up to 60% liberal or conservative in samples as small as $n_{1}=n_{2}=10$. It is also of interest to note that, Dudewicz et al. (2007) constructed exact tests, but only supply critical constants—not a symbolic density. Their method does not reveal the algebraic structure (hypergeometric kernel, termination when $\nu_{i}$ is odd, Ramanujan tails) that our MB analysis uncovers. Their exactness is achieved by numerical double integration for every $(n_{1},n_{2},r)$. We give an expression that works for all $(n_{1},n_{2},r)$ in a single expression.
The novelty of our new approach is that it introduces a unified analytic mechanism for handling the core obstruction appearing in the Behrens–Fisher statistic: the coupled term in the denominator involving a weighted sum of two independent chi-square variates. The classical literature on the Behrens–Fisher problem is dominated by three principal lines of development: (i) approximation-based methods (most notably, Welch (1947), Satterthwaite (1946)), (ii) exact or near-exact testing procedures based on numerical integration or generalized pivotal quantities, and (iii) special-case representations in terms of hypergeometric functions. What has been missing is a systematic framework that explains—in a unified manner—why a hypergeometric structure emerges, why particular finite-sum representations occur in special cases, and how one can transition from the central (bulk) region of the distribution to its tail behavior without modifying the underlying analytic framework. The MB contour representation accomplishes precisely this task, thereby distinguishing our work from the existing literature: decomposing the troublesome square-root factor into a contour integral converts the original two-dimensional convolution into a one-dimensional object whose singularities are completely explicit.
Earlier work, including the important result of Nel et al. (1990), identified a hypergeometric pdf, but did not develop the underlying MB contour decomposition structure that produces it. In our treatment, the hypergeometric form is not an isolated formula; rather, it arises as one manifestation of a broader analytic framework. The residue expansion explains the terminating algebraic cases; the Euler–Beta reduction yields the same density derived by Nel et al. (1990) in compact hypergeometric form and extends naturally to the cdf; RMT provides explicit tail coefficients; and the same cumulant structure underpins an LR saddle-point approximation with quantified error. In essence, our contribution ties together the exact density theory, finite-sum special cases, tail asymptotics, and numerical evaluation in a manner that complements and extends existing treatments for the Behrens–Fisher problem.
@@
The remainder of the paper is organized as follows. In Section 2, the MB representation of $T$ is presented. The density function $f_{T}(t)$ is expressed as a Gauss hypergeometric function $_{2}F_{1}$ by reducing the 2-D integral to a 1-D integral in Section 3. The tail analyses of $f_{T}(t)$ using Ramanujan’s master theorem is in Section 4. In Section 5, we provide discussions including specifications for cdf tables, LR saddle-points, and comparisons between our work with Welch (1947) and Nel et al. (1990). Section 6 concludes the paper.